solve for x,y, and z in the system of equations below. Explain your steps!!!???
x+7y+3z=29
3z+x-2y=-7
5y=10-2x

Answer by TimothyLamb(4379)

---
x+7y+3z=29
3z+x-2y=-7
5y=10-2x
---
put the system of linear equations into standard form:
---
x+7y+3z=29
x-2y+3z=-7
2x+5y=10
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution by matrix method:
x = -5
y = 4
z = 2
---
now elimination method:
---
x+7y+3z=29
(x-2y+3z=-7)*-1
---
x+7y+3z=29
-x+2y-3z=7
---
add subsystem:
---
9y=36
y = 4
---
(x+7y+3z=29)*-2
2x+5y=10
---
-2x-14(4)-6z=-58
2x+5(4)=10
---
add subsystem:
---
-36 - 6z =-48
-6z =-12
z = 2
---
x+7y+3z=29
x=29-7y-3z
x=29-7(4)-3(2)
x = -5
---
solution by elimination method:
x = -5
y = 4
z = 2
---
both methods agree !
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php